OpenStudy (anonymous):

Please help me finish my super long hw. A candy store makes an 11-pound mixture of gummy candy, jelly beans, and hard candy. The cost of gummy candy is $2.00 per pound, jelly beans cost $3.00 per pound, and hard candy costs $3.00 per pound. The mixture calls for two times as many gummy candy pieces as jelly beans. The total cost of the mixture is $27.00. How much of each ingredient did the store use?

OpenStudy (anonymous):

maybe try averaging I don't really know!

OpenStudy (anonymous):

idk, idr how i was taught to do this but my teacher isn't in today so i can't ask him for help :(

OpenStudy (anonymous):

im doing online school and registered late so he bumped me up a unit so I could catch up and im taking a test and I need help right now! haha!

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

Well, if i get it right: For each pound of Jelly beans there have to be 2 pounds of Gummy candy, right? So that means we can combine them like so: \[ mix = 1 \cdot Jelly + 2 \cdot Gummy = $3 + 2 \cdot $2 = $7 \] Notice that mix is 3 pounds, 1 for each ingredient in it. Now, $7 could get 3 times in $27 That means $7 * 3 = $21 and also 3pounds * 3 = 9 pounds Now to get to $27 all you have to do is add another 2 pounds of the Hard candies which is $3 each.. And you get a $27 mixture that is 11 pounds. Is that good enough? Does it have to be more eh... formal?